1) Suppose the marginal product of labor in the economy is given by MPN = 0.005(18,000 - .005N), while the supply of labor is 2000 + 1000w. (a) Find the market-clearing real wage rate and level of employment. (b) What happens to the wage rate and employment if wealth rises, reducing the supply of labor to 200 + 1000w?

2) Find the interest rate that clears the goods market of an economy that has has full-employment output of 5000. Government purchases are 1000. Desired consumption and desired investment are given by Cd = 3000 - 2000r + 0.10Y I d = 1000 - 4000r where Y is output and r is the expected real interest rate. Given your clearing real interest rate what is the amount of saving, investment, and consumption in equilibrium.
Answer #1

Solution

(a) Market Clearing real wage rate and level of employment :-

The market-clearing real wage rate equates the demand and supply of labor.

Setting w = MPN

w = 90 - 0.005(2000+1000w)

w = 90-10 - 5w

w = 80-5w

6w = 80

w = 13.33

The wage rate will be 13.33

Putting this in labor supply equation,

N=2000 + 1000 × 13.33

N=15,330

The employment will be 15,330

(b)

Setting w = MPN = 0.005(18,000 - .005N)

w = 90 -0.005(200 + 1000w)

w = 90-1 - 5w

6w = 89

w = 14.83

The wage rate increased to 14.83

Putting this in labor supply,

N = 200 + 1000 × 14.83 = 15030

The employment will be reduced to 15030